OpenStudy (anonymous):

Find the LCM (7+3x),(49-9x^2),(7-3x)

OpenStudy (anonymous):

notice that the middle expression is the product of the other two: (7+3x)*(7-3x) = 49-9x^2 so your answer should be...

OpenStudy (ash2326):

We have \[(7+3x), (49-9x^2), (7-3x)\] LCM = \[(the\ common\ factor\ \times the\ uncommon\ factors)\] Let's factor\( (49-9x^2)\) We know that \[a^2-b^2=(a+b)(a-b)\] so \[49-9x^2=(7+3x)(7-3x)\] so We have \[7+3x, (7+3x)(7-3x), (7-3x)\] Let's find the LCM of the first two terms Common Factor= (7+3x) Uncommon Factor for first = none Uncommon factor for second term= (7-3x) So LCM is \[(7+3x) \times(7-3x)\] Now take LCM of \((7+3x) \times(7-3x)\) and (7-3x) Common factor = 7-3x Uncommon for first= 7+3x Uncommon for second = none so \[LCM\ is= (7+3x)(7-3x)\]

OpenStudy (anonymous):

impressive... try thinking about your problem this way... What's the LCM of 2,6,3? Since 2*3=6, LCM must be 6.

OpenStudy (anonymous):

Which is equal to 21x(x+3)(x+7)